OpenStudy (anonymous):

please show that the sequence a(n+1)=sqr(3an) where a1=2 is convergent and find its limit.

OpenStudy (alekos):

do you still need help on this?

OpenStudy (anonymous):

To show the sequence converges, you have to establish monotonicity and boundedness. To see if there's an initial pattern of monotonicity, compare the first two terms: \[a_2=\sqrt{3a_1}=\sqrt{3\times2}=\sqrt6>\sqrt4=2=a_1\] So we know the sequence is increasing at least between the first two terms. This is the basis case (\(n=1\)). Assume \(a_{k+1}>a_k\), and show that \(a_{k+2}>a_{k+1}\). \[\begin{align*} a_{k+2}&=\sqrt{3a_{k+1}}\\ &>\sqrt{3a_k}\\ &=a_{k+1} \end{align*}\] and so the sequence is increasing for all \(n\ge1\). To establish boundedness, you can notice as sort of pattern: \[\begin{array}{cr}a_1&&2<3\\ a_2&&\sqrt6<\sqrt9=3\\ a_3&&\sqrt{3\sqrt6}=\sqrt[4]{54}<\sqrt[4]{81}=3\\ a_4&&\sqrt{3\sqrt[4]{54}}=\sqrt[8]{4374}<\sqrt[8]{6561}=3 \end{array}\] and so on. So we know that 3 is an upper bound. The sequence is bounded and monotonic, so it must converge.